# Online Smoothed Demand Management

### v1

@@ 8 Conclusion
Motivated by paradigm shifts promoting active and intelligent demand management models (e.g., demand response), we introduce and study online smoothed demand management (OSDM) – the first online problem to jointly capture key practical features such as mixed inflexible and flexible demand, local energy storage, and costs penalizing unsmooth behavior. We present PAAD, an algorithm achieving the optimal competitive ratio for OSDM, and PALD, a novel differentiable framework that learns the best algorithm from data while preserving worst-case guarantees. We evaluate both in a case study of a grid-integrated data center with co-located storage. Several directions remain for future work. One natural extension is to incorporate rate constraints, where purchase and delivery rates are capped per time step – a constraint captured by related models Lechowicz:24; Yang:20. More broadly, while our differentiable framework is developed for OSDM, its end-to-end learning approach may generalize to other online problems, motivating further study of which problem classes and algorithm parameterizations admit such robust learning.
Acknowledgments
This research is supported by National Science Foundation grants 2045641, 2325956, 2512128, and 2533814, and the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0024386.
Disclaimers
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
@@ Appendix A Deferred Experiment Setup and Results
In this section, we provide additional experimental results and details that were deferred from § 7. We start by discussing the runtime of our algorithms in § A.1, before presenting additional experimental results in § A.2 and presenting details of the PALD implementations in § A.3. Note that Fig. 9 plots the “weighted active jobs” metric used to model demand using the Alibaba trace [Alibaba:18], described in § 7.
Figure 6: Weighted active jobs over time (bucketed in 15-min intervals) for the Alibaba trace [Alibaba:18]
Figure 7: Average per-time-step runtime (in ms) for all algorithms with varying instance lengths $T$.
Figure 8: Average ECR for all algorithms in CAISO with varying delivery switching coefficients $\delta$.
Figure 9: Average ECR for PALD-S in default CAISO experiment trained with varying $\rho$.
A.1 Runtime Measurements, Discussion, and Optimizations
@@
In Fig. 9, we show the average per-time-step runtime (in milliseconds) for all algorithms with varying instance lengths $T$, averaging over 8,400 instances (1,200 for each setting of $T$). We also include the runtime of the offline solver (computed as a bilinear program using GurobiPy [gurobi]), normalized by the instance length $T$ to give a fair comparison with the online algorithms. Each instance and each algorithm was run in a single thread on a MacBook Air with M2 processor and 24 GB of RAM. Unsurprisingly, we find that the offline optimal solution is the most computationally expensive, exhibiting a “per-time-step” runtime an order of magnitude higher than the next closest online algorithm. PAAD and PALD-C have comparable runtimes, with averages of 62.5 ms and 65.7 ms per time step, respectively, which remains practical for real-time decision-making.
@@
PALD-S’s performance improvement demonstrates that several optimizations can be used to speed-up the core psuedo-cost minimization problem, e.g., by caching results, using approximations, or re-parameterizing the integral – we do not explore these optimizations in depth in this work, but they offer further evidence to suggest that the PAAD and PALD frameworks are practical for real-time decision-making in large-scale systems.
We also remark that additional driver-level optimizations can be employed to reduce the number of times that the psuedo-cost problem needs to be solved without affecting the theoretical results. Perhaps the highest impact of these optimizations is driver consolidation – if multiple drivers have the same type and relative state, they can be consolidated into a single driver with a larger size, allowing one solution to the pseudo-cost problem to essentially cover multiple drivers. Formally, if for example, two drivers $i$ and $j$ have the same type (i.e., $d^{(i)}=d^{(j)}$) and the same relative state (i.e., $\nicefrac{{w_{b}^{(i)}}}{{d^{(i)}}}=\nicefrac{{w_{b}^{(j)}}}{{d^{(j)}}}$), then they can be consolidated into a single driver with size $d^{(i)}+d^{(j)}$. We did not implement this optimization in our experiments, but it could be effective in practice to even further reduce the number of times that the pseudo-cost problem needs to be solved.
A.2 Deferred Experimental Results
@@
$\triangleright$ Delivery Switching Cost. Fig. 9 shows the effect of varying the switching cost coefficient $\delta$ (with $\gamma$ fixed at $\gamma=10$) between 0 and $12.5$ in CAISO – other parameters are set to the experiment defaults (see § 7). We find that the performance of all algorithms slightly improves as the switching cost coefficient $\delta$ increases.
@@
A.3 PALD Implementation Details
@@
We implement the PALD algorithm using CVXPYLayers [CVXPYlayers:19, CVXPY] to enable optimization of the downstream task (i.e., minimizing the empirical competitive ratio on training instances) via backpropagation through the optimization problems that define PALD’s decisions.
As outlined in § 6.2, PALD-S directly learns one set of piecewise-affine threshold functions to minimize overall empirical competitive ratio on training instances. We implement projected gradient descent using PyTorch [Paszke:19:PyTorch] and CVXPY [CVXPY] to project the learned thresholds (specifically the parameter vectors) into the feasible sets. To capture some degree of seasonality, we train a set of PALD-S thresholds for each month in a given region’s LMP data (i.e., 12 sets of thresholds in total), using 100 random training instances for each month. In Fig. 10, we give a high-level diagram of the PALD-S learning framework.
Figure 10: PALD-S learns a single set of thresholds (via the parameter vectors $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi})$ to directly optimize a downstream task loss of the empirical competitive ratio.
@@
Figure 11: PALD-C learns a mapping between context about each driver (see Algorithm 2) and the instance-specific threshold functions that minimize overall task loss. Gradients flow through the differentiable PALD implementation and the final projection layer of the neural network to improve this mapping.
In our implementation of both PALD-S and PALD-C, we use $K=10$ grid points to parameterize the piecewise-affine threshold functions. We restrict the thresholds to lie in a concave region by enforcing that the second-order finite differences of the parameter vectors are non-positive. The reason we do this is because we use a nonnegative sum of squared hinges to parameterize the computation of the integral over the threshold functions in the pseudo-cost minimization problem (i.e., $\int_{w}^{w+x}\phi_{b}(u)du$, $\int_{w}^{w+x}\phi_{f}(u)du$, and $\int_{v}^{v+x}\psi_{f}(u)du$, see Algorithm 1 for details). This parameterization is convex and DPP (s disciplined parameterized program) in $x$ if the thresholds are linear decreasing or concave decreasing. We implement this parameterization to employ CVXPYLayers, which require the optimization problem to be DPP in the optimization variables (i.e., $x$) and the parameters (i.e., the threshold function parameters). However, we note that this concavity constraint is not required for the theoretical guarantees of PALD and could likely be circumvented with a different implementation – the underlying integral is analytically convex in $x$ for any non-increasing threshold function.
@@ Appendix B Deferred Discussion from § 2
Grid-integrated Data Center with Storage. Consider a grid-connected data center with local energy storage. At each discrete time step $t$ (e.g., 15-minute intervals), the data center must decide how much energy to purchase ($x_{t}$) from the grid at a price $p_{t}$, and how much energy to deliver (i.e., use now) to meet demand ($z_{t}$) – the difference between these quantities is stored in (resp. delivered from) the local energy storage with state of charge $s_{t}$. In this setting, base demands $b_{t}$ model e.g., interactive workloads that must be satisfied immediately, while flexible demands $f_{t}$ model e.g., delay-tolerant batch jobs that can be deferred until a deadline $\Delta_{t}>t$. The data center’s goal is to minimize its electricity cost ($p_{t}x_{t}$) while smoothing its purchasing rate ($\mathcal{S}(\cdot)$), since large fluctuations in electricity consumption negatively impact the grid [Bloomberg:24]. The data center may also wish to place a small penalty on the cost of delivering energy ($\mathcal{D}(\cdot)$) from the battery to e.g., amortize the battery’s degradation over time, as well as a penalty on the smoothness of the delivery rate (i.e., its own rate of energy consumption within the data center) to reduce wear-and-tear on internal components ($\delta|z_{t}-z_{t-1}|$). Compared to prior work that has studied demand management in data centers [Yang:20], OSDM captures the practical considerations of both flexible and inflexible demand, as well as the desire to make “smooth” decisions on the grid- and demand-sides of the problem.
@@
Thermal Energy Demand Management. Consider a thermal energy system with a local energy storage tank [NREL:ThermalStorage]. At each time step $t$, the operator must decide how much thermal energy to purchase ($x_{t}$) at a price $p_{t}$ (e.g., by converting grid electricity into heat), and how much thermal energy to deliver to meet demand ($z_{t}$) – the difference between these quantities is stored in (resp. delivered from) the local storage tank with state of charge $s_{t}$. Base demands $b_{t}$ model e.g., heating or cooling loads that must be satisfied immediately, while flexible demands $f_{t}$ model e.g., delay-tolerant heating or cooling loads that can be deferred until a deadline $\Delta_{t}>t$. The operator’s goal is to minimize their energy cost ($p_{t}x_{t}$) while smoothing their purchasing rate ($\mathcal{S}(\cdot)$). The delivery cost ($\mathcal{D}(\cdot)$) captures the extra input energy required to drive the delivery system (e.g., pumps, heat exchangers, etc.), which may depend on the price of energy ($p_{t}$). The operator may also wish to place a small penalty on the smoothness of the delivery rate (i.e., its own rate of thermal energy consumption) to reduce wear-and-tear on distribution components such as valves ($\delta|z_{t}-z_{t-1}|$).
@@
Just-in-Time Manufacturing with Material Inventory. Consider a factory that processes raw materials into a finished product, and has a warehouse to store raw materials. At each time step $t$, the factory must decide how much raw material to purchase ($x_{t}$) at a price $p_{t}$, and how much finished product to deliver to meet demand ($z_{t}$) – the difference between these quantities is stored in (resp. delivered from) the warehouse with state $s_{t}$. Orders for the product arrive over time, where base demands $b_{t}$ model e.g., orders that must be satisfied immediately, while flexible demands $f_{t}$ model scheduled orders that can be deferred until a deadline $\Delta_{t}>t$. The factory’s goal is to minimize their material cost ($p_{t}x_{t}$) while smoothing their purchasing rate ($\mathcal{S}(\cdot)$), which can reduce supply chain uncertainty and lead to discounts from suppliers. The delivery cost ($\mathcal{D}(\cdot)$) captures the processing cost for converting raw materials into the finished product at time $t$. The factory may also wish to place a small penalty on the smoothness of the delivery rate (i.e., its own rate of production) to encourage a more predictable production schedule ($\delta|z_{t}-z_{t-1}|$).
@@
Flow Battery Storage Management. Consider a large energy consumer that operates a grid-connected flow battery to store energy during off-peak hours. Flow batteries are a type of rechargeable battery where energy is stored in the form of two (or more) liquids contained in external tanks, and energy is extracted by pumping the liquids through a cell stack to generate electricity [Breeze:19]. At each time step $t$, the operator must decide how much electricity to purchase ($x_{t}$) from the grid at a price $p_{t}$, and how much electricity to deliver to meet demand ($z_{t}$) – the difference between these quantities is stored in (resp. delivered from) the flow battery with state of charge $s_{t}$. The operator’s goal is to minimize their electricity cost ($p_{t}x_{t}$) while smoothing their purchasing rate ($\mathcal{S}(\cdot)$). The delivery cost ($\mathcal{D}(\cdot)$) captures the extra input energy required to drive the flow battery system (e.g., pumps, valves, etc.), which may depend on the price of energy ($p_{t}$).
@@ Appendix C Details of the OCS Problem and the RORO Algorithm
In this section, we provide deferred details about the OCS problem and the “ramp-on, ramp-off” (RORO) algorithm proposed to solve it optimally [Lechowicz:24], which we discuss in the warmup (see § 3). We start by formally defining the OCS problem, which can be cast as a special case of OSDM with a single unit of flexible demand and no storage. We then describe the RORO algorithm and its guarantee in § C.2.
C.1 Online Conversion with Switching Costs (OCS) [Lechowicz:24]
@@
$[\texttt{OCS}]\quad\min_{\mathbf{x}\coloneqq\{x_{t}\}_{t\in T}}$ $\underbrace{\sum_{t=1}^{T}p_{t}x_{t}}_{\text{total purchasing cost}}+\underbrace{\sum_{t=1}^{T+1}\gamma|x_{t}-x_{t-1}|,}_{\text{switching penalty}}$ (5)
s.t. $\underbrace{\sum_{t=1}^{T}x_{t}=1,}_{\text{deadline constraint}}\quad x_{t}\in[0,d_{t}]\ \forall t\in[T].$ (6)
@@
Lechowicz:24 focus on the online version of OCS, where the decision-maker can only observe price signals up to time $t$ when selecting $x_{t}$, and each choice of $x_{t}$ is irrevocable (i.e., it cannot be revised at future time steps). The goal is to design an online algorithm that achieves a small competitive ratio (see Def. 2.1) They make the following assumptions on the problem:
@@
Assumptions. Prices have bounded support, i.e., $p_{t}\in[p_{\min},p_{\max}]\ \forall t\in[T]$, where $p_{\min}$ and $p_{\max}$ are known positive constants. They also assume that the switching coefficient $\gamma$ is “not too large”. Formally, it is bounded within $\gamma\in\left[0,\frac{(p_{\max}-p_{\min})}{2}\right)$. If $\gamma$ exceeds this range, its impact on the total cost exceeds that of the prices, and any competitive algorithm should simply minimize this term, making decision-making trivial [Lechowicz:24]. They typically assume that the deadline $T$ is known in advance to facilitate a “compulsory trade” that ensures the entire asset is purchased before the end of the sequence. If the operator has completed $w_{T}$ fraction of the workload at time $T$, they are forced to purchase the remaining $1-w_{T}$ fraction at time $T$.
C.2 The RORO Algorithm [Lechowicz:24]
@@
In the classic competitive setting, OCS is solved optimally by an online algorithm framework called “ramp-on, ramp-off” (RORO), shown by [Lechowicz:24]. In what follows, we present this algorithm to contextualize our results in § 3.
@@
In RORO [Lechowicz:24], the online decision at each time step is made by solving a pseudo-cost minimization problem to determine the amount to purchase at the current time step (i.e., $x_{t}\in[0,1-w_{t-1}]$). This minimization balances between the extreme options of buying “too much” early (thus incurring suboptimal costs if prices later drop) and waiting too long to purchase ( risking being forced to purchase a large portion at once, potentially at high cost).
@@
Whenever the price is “sufficiently attractive”, the pseudo-cost minimization finds the best decision that purchase just enough to maintain a certain competitive ratio. To define this trade-off, Lechowicz:24 introduce a dynamic threshold function $\phi(w):[0,1]\rightarrow[p_{\min},p_{\max}]$, which is a function that maps the current amount that has been purchased $w$ to a price value. It is defined as:
$\phi(w)=p_{\max}-\gamma+\left(\frac{p_{\max}}{\alpha_{\texttt{RORO}}}-p_{\max}+2\gamma\right)\exp\left(\frac{w}{\alpha_{\texttt{RORO}}}\right),$ (7)
@@
where $\alpha_{\texttt{RORO}}$ is the optimal competitive ratio for OCS, defined as the unique positive solution to $\exp\left(\frac{1}{\alpha_{\texttt{RORO}}}\right)=\frac{p_{\max}-p_{\min}-2\gamma}{\nicefrac{{p_{\max}}}{{\alpha_{\texttt{RORO}}}}-p_{\max}-2\gamma}$ [Lechowicz:24]:
@@
$\alpha_{\texttt{RORO}}\coloneqq\left[W\left(\left(\frac{2\gamma+p_{\min}}{p_{\max}}-1\right)\exp\left(\frac{2\gamma}{p_{\max}}-1\right)\right)-\frac{2\gamma}{p_{\max}}+1\right]^{-1}.$ (8)
@@
In the above, $W(\cdot)$ is the Lambert $W$ function [Corless:96LambertW]. Given this definition of $\alpha_{\texttt{RORO}}$, note that $\phi(\cdot)$ is monotonically decreasing on the interval $w\in[0,1]$ (i.e., non-increasing in the amount of demand that has already been purchased), and captures the marginal trade-off of making progress towards satisfying the single unit of demand. We summarize the RORO algorithm in Algorithm 3.
Algorithm 3 Online Ramp-On, Ramp-Off Algorithm (RORO[Lechowicz:24]

### v2

@@ 8. Conclusion
Motivated by paradigm shifts promoting active and intelligent demand management models (e.g., demand response), we introduce and study online smoothed demand management (OSDM)—the first online problem to jointly capture key practical features such as mixed inflexible and flexible demand, local energy storage, and costs penalizing unsmooth behavior. We present PAAD, an algorithm achieving the optimal competitive ratio for OSDM, and PALD, a novel differentiable framework that learns the best algorithm from data while preserving worst-case guarantees. We evaluate both in a case study of a grid-integrated data center with co-located storage.
Several directions remain for future work. One natural extension is to incorporate rate constraints, where purchase and/or delivery rates are capped per time step as in some related models (Lechowicz et al., 2024; Yang et al., 2020). While the monotone state- and price-dependent delivery cost we consider (Def. 2.4) is sufficiently general to capture some motivating applications, the delivery cost could be generalized—for instance, one might consider a cost that depends on a second adversarial price sequence (separate from $p_{t}$). More broadly, while our PALD learning framework is developed for OSDM, its approach of learning “the best algorithm” out of a set of algorithms that satisfy a given worst-case competitive ratio may generalize to other online problems, motivating further study of which problem classes and algorithm parameterizations admit such a framework.
@@ Appendix A Deferred Experiment Setup and Results
In this section, we provide additional experimental results and details that were deferred from § 7. We start by discussing the runtime of our algorithms in § A.1, before presenting additional experimental results in § A.2 and presenting details of the PALD implementations in § A.3. Note that Fig. 9 plots the “weighted active jobs” metric used to model demand using the Alibaba trace (Alibaba, 2018), described in § 7.
Figure 6. Weighted active jobs over time (bucketed in 15-min intervals) for the Alibaba trace (Alibaba, 2018)
Figure 7. Average per-time-step runtime (in ms) for all algorithms with varying instance lengths $T$.
Figure 8. Average ECR for all algorithms in CAISO with varying delivery switching coefficients $\delta$.
Figure 9. Average ECR for PALD-S in default CAISO experiment trained with varying $\rho$.
A.1. Runtime Measurements, Discussion, and Optimizations
@@
In Fig. 9, we show the average per-time-step runtime (in milliseconds) for all algorithms with varying instance lengths $T$, averaging over 8,400 instances (1,200 for each setting of $T$). We also include the runtime of the offline solver (computed as a bilinear program using GurobiPy (Gurobi Optimization, LLC, 2024)), normalized by the instance length $T$ to give a fair comparison with the online algorithms. Each instance and each algorithm was run in a single thread on a MacBook Air with M2 processor and 24 GB of RAM. Unsurprisingly, we find that the offline optimal solution is the most computationally expensive, exhibiting a “per-time-step” runtime an order of magnitude higher than the next closest online algorithm. PAAD and PALD-C have comparable runtimes, with averages of 62.5 ms and 65.7 ms per time step, respectively, which remains practical for real-time decision-making.
@@
PALD-S’s performance improvement demonstrates that several optimizations can be used to speed-up the core psuedo-cost minimization problem, e.g., by caching results, using approximations, or re-parameterizing the integral—we do not explore these optimizations in depth in this work, but they offer further evidence to suggest that the PAAD and PALD frameworks are practical for real-time decision-making in large-scale systems.
We also remark that additional driver-level optimizations can be employed to reduce the number of times that the psuedo-cost problem needs to be solved without affecting the theoretical results. Perhaps the highest impact of these optimizations is driver consolidation—if multiple drivers have the same type and relative state, they can be consolidated into a single driver with a larger size, allowing one solution to the pseudo-cost problem to essentially cover multiple drivers. Formally, if for example, two drivers $i$ and $j$ have the same type (i.e., $d^{(i)}=d^{(j)}$) and the same relative state (i.e., $\nicefrac{{w_{b}^{(i)}}}{{d^{(i)}}}=\nicefrac{{w_{b}^{(j)}}}{{d^{(j)}}}$), then they can be consolidated into a single driver with size $d^{(i)}+d^{(j)}$. We did not implement this optimization in our experiments, but it could be effective in practice to even further reduce the number of times that the pseudo-cost problem needs to be solved.
A.2. Deferred Experimental Results
@@
$\triangleright$ Delivery Switching Cost. Fig. 9 shows the effect of varying the switching cost coefficient $\delta$ (with $\gamma$ fixed at $\gamma=10$) between 0 and $12.5$ in CAISO—other parameters are set to the experiment defaults (see § 7). We find that the performance of all algorithms slightly improves as the switching cost coefficient $\delta$ increases.
@@
A.3. PALD Implementation Details
@@
We implement the PALD algorithm using CVXPYLayers (Agrawal et al., 2019; Diamond and Boyd, 2016) to enable optimization of the downstream task (i.e., minimizing the empirical competitive ratio on training instances) via backpropagation through the optimization problems that define PALD’s decisions.
As outlined in § 6.2, PALD-S directly learns one set of piecewise-affine threshold functions to minimize overall empirical competitive ratio on training instances. We implement projected gradient descent using PyTorch (Paszke et al., 2019) and CVXPY (Diamond and Boyd, 2016) to project the learned thresholds (specifically the parameter vectors) into the feasible sets. To capture some degree of seasonality, we train a set of PALD-S thresholds for each month in a given region’s LMP data (i.e., 12 sets of thresholds in total), using 100 random training instances for each month. In Fig. 10, we give a high-level diagram of the PALD-S learning framework.
Figure 10. PALD-S learns a single set of thresholds (via the parameter vectors $\mathbf{y}_{b},\mathbf{y}_{f},\mathbf{y}_{\psi})$ to directly optimize a downstream task loss of the empirical competitive ratio.
@@
Figure 11. PALD-C learns a mapping between context about each driver (see Algorithm 2) and the instance-specific threshold functions that minimize overall task loss. Gradients flow through the differentiable PALD implementation and the final projection layer of the neural network to improve this mapping.
In our implementation of both PALD-S and PALD-C, we use $K=10$ grid points to parameterize the piecewise-affine threshold functions. We restrict the thresholds to lie in a concave region by enforcing that the second-order finite differences of the parameter vectors are non-positive. The reason we do this is because we use a nonnegative sum of squared hinges to parameterize the computation of the integral over the threshold functions in the pseudo-cost minimization problem (i.e., $\int_{w}^{w+x}\phi_{b}(u)du$, $\int_{w}^{w+x}\phi_{f}(u)du$, and $\int_{v}^{v+x}\psi_{f}(u)du$, see Algorithm 1 for details). This parameterization is convex and DPP (s disciplined parameterized program) in $x$ if the thresholds are linear decreasing or concave decreasing. We implement this parameterization to employ CVXPYLayers, which require the optimization problem to be DPP in the optimization variables (i.e., $x$) and the parameters (i.e., the threshold function parameters). However, we note that this concavity constraint is not required for the theoretical guarantees of PALD and could likely be circumvented with a different implementation—the underlying integral is analytically convex in $x$ for any non-increasing threshold function.
@@ Appendix B Deferred Discussion from § 2
Grid-integrated Data Center with Storage. Consider a grid-connected data center with local energy storage. At each discrete time step $t$ (e.g., 15-minute intervals), the data center must decide how much energy to purchase ($x_{t}$) from the grid at a price $p_{t}$, and how much energy to deliver (i.e., use now) to meet demand ($z_{t}$)—the difference between these quantities is stored in (resp. delivered from) the local energy storage with state of charge $s_{t}$. In this setting, base demands $b_{t}$ model e.g., interactive workloads that must be satisfied immediately, while flexible demands $f_{t}$ model e.g., delay-tolerant batch jobs that can be deferred until a deadline $\Delta_{t}>t$. The data center’s goal is to minimize its electricity cost ($p_{t}x_{t}$) while smoothing its purchasing rate ($\mathcal{S}(\cdot)$), since large fluctuations in electricity consumption negatively impact the grid (Nicoletti et al., 2024). The data center may also wish to place a small penalty on the cost of delivering energy ($\mathcal{D}(\cdot)$) from the battery to e.g., amortize the battery’s degradation over time, as well as a penalty on the smoothness of the delivery rate (i.e., its own rate of energy consumption within the data center) to reduce wear-and-tear on internal components ($\delta|z_{t}-z_{t-1}|$). Compared to prior work that has studied demand management in data centers (Yang et al., 2020), OSDM captures the practical considerations of both flexible and inflexible demand, as well as the desire to make “smooth” decisions on the grid- and demand-sides of the problem.
@@
Thermal Energy Demand Management. Consider a thermal energy system with a local energy storage tank (National Renewable Energy Laboratory, 2024). At each time step $t$, the operator must decide how much thermal energy to purchase ($x_{t}$) at a price $p_{t}$ (e.g., by converting grid electricity into heat), and how much thermal energy to deliver to meet demand ($z_{t}$)—the difference between these quantities is stored in (resp. delivered from) the local storage tank with state of charge $s_{t}$. Base demands $b_{t}$ model e.g., heating or cooling loads that must be satisfied immediately, while flexible demands $f_{t}$ model e.g., delay-tolerant heating or cooling loads that can be deferred until a deadline $\Delta_{t}>t$. The operator’s goal is to minimize their energy cost ($p_{t}x_{t}$) while smoothing their purchasing rate ($\mathcal{S}(\cdot)$). The delivery cost ($\mathcal{D}(\cdot)$) captures the extra input energy required to drive the delivery system (e.g., pumps, heat exchangers, etc.), which may depend on the price of energy ($p_{t}$). The operator may also wish to place a small penalty on the smoothness of the delivery rate (i.e., its own rate of thermal energy consumption) to reduce wear-and-tear on distribution components such as valves ($\delta|z_{t}-z_{t-1}|$).
@@
Just-in-Time Manufacturing with Material Inventory. Consider a factory that processes raw materials into a finished product, and has a warehouse to store raw materials. At each time step $t$, the factory must decide how much raw material to purchase ($x_{t}$) at a price $p_{t}$, and how much finished product to deliver to meet demand ($z_{t}$)—the difference between these quantities is stored in (resp. delivered from) the warehouse with state $s_{t}$. Orders for the product arrive over time, where base demands $b_{t}$ model e.g., orders that must be satisfied immediately, while flexible demands $f_{t}$ model scheduled orders that can be deferred until a deadline $\Delta_{t}>t$. The factory’s goal is to minimize their material cost ($p_{t}x_{t}$) while smoothing their purchasing rate ($\mathcal{S}(\cdot)$), which can reduce supply chain uncertainty and lead to discounts from suppliers. The delivery cost ($\mathcal{D}(\cdot)$) captures the processing cost for converting raw materials into the finished product at time $t$. The factory may also wish to place a small penalty on the smoothness of the delivery rate (i.e., its own rate of production) to encourage a more predictable production schedule ($\delta|z_{t}-z_{t-1}|$).
@@
Flow Battery Storage Management. Consider a large energy consumer that operates a grid-connected flow battery to store energy during off-peak hours. Flow batteries are a type of rechargeable battery where energy is stored in the form of two (or more) liquids contained in external tanks, and energy is extracted by pumping the liquids through a cell stack to generate electricity (Breeze, 2019). At each time step $t$, the operator must decide how much electricity to purchase ($x_{t}$) from the grid at a price $p_{t}$, and how much electricity to deliver to meet demand ($z_{t}$)—the difference between these quantities is stored in (resp. delivered from) the flow battery with state of charge $s_{t}$. The operator’s goal is to minimize their electricity cost ($p_{t}x_{t}$) while smoothing their purchasing rate ($\mathcal{S}(\cdot)$). The delivery cost ($\mathcal{D}(\cdot)$) captures the extra input energy required to drive the flow battery system (e.g., pumps, valves, etc.), which may depend on the price of energy ($p_{t}$).
@@ Appendix C Details of the OCS Problem and the RORO Algorithm
In this section, we provide deferred details about the OCS problem and the “ramp-on, ramp-off” (RORO) algorithm proposed to solve it optimally (Lechowicz et al., 2024), which we discuss in the warmup (see § 3). We start by formally defining the OCS problem, which can be cast as a special case of OSDM with a single unit of flexible demand and no storage. We then describe the RORO algorithm and its guarantee in § C.2.
C.1. Online Conversion with Switching Costs (OCS) (Lechowicz et al., 2024)
@@
(6) $[\texttt{OCS}]\quad\min_{\mathbf{x}\coloneqq\{x_{t}\}_{t\in T}}$ $\underbrace{\sum_{t=1}^{T}p_{t}x_{t}}_{\text{total purchasing cost}}+\underbrace{\sum_{t=1}^{T+1}\gamma|x_{t}-x_{t-1}|,}_{\text{switching penalty}}$
(7) s.t. $\underbrace{\sum_{t=1}^{T}x_{t}=1,}_{\text{deadline constraint}}\quad x_{t}\in[0,d_{t}]\ \forall t\in[T].$
@@
Lechowicz et al. (2024) focus on the online version of OCS, where the decision-maker can only observe price signals up to time $t$ when selecting $x_{t}$, and each choice of $x_{t}$ is irrevocable (i.e., it cannot be revised at future time steps). The goal is to design an online algorithm that achieves a small competitive ratio (see Def. 2.1) They make the following assumptions on the problem:
@@
Assumptions. Prices have bounded support, i.e., $p_{t}\in[p_{\min},p_{\max}]\ \forall t\in[T]$, where $p_{\min}$ and $p_{\max}$ are known positive constants. They also assume that the switching coefficient $\gamma$ is “not too large”. Formally, it is bounded within $\gamma\in\left[0,\frac{(p_{\max}-p_{\min})}{2}\right)$. If $\gamma$ exceeds this range, its impact on the total cost exceeds that of the prices, and any competitive algorithm should simply minimize this term, making decision-making trivial (Lechowicz et al., 2024). They typically assume that the deadline $T$ is known in advance to facilitate a “compulsory trade” that ensures the entire asset is purchased before the end of the sequence. If the operator has completed $w_{T}$ fraction of the workload at time $T$, they are forced to purchase the remaining $1-w_{T}$ fraction at time $T$.
C.2. The RORO Algorithm (Lechowicz et al., 2024)
@@
In the classic competitive setting, OCS is solved optimally by an online algorithm framework called “ramp-on, ramp-off” (RORO), shown by (Lechowicz et al., 2024). In what follows, we present this algorithm to contextualize our results in § 3.
@@
In RORO (Lechowicz et al., 2024), the online decision at each time step is made by solving a pseudo-cost minimization problem to determine the amount to purchase at the current time step (i.e., $x_{t}\in[0,1-w_{t-1}]$). This minimization balances between the extreme options of buying “too much” early (thus incurring suboptimal costs if prices later drop) and waiting too long to purchase ( risking being forced to purchase a large portion at once, potentially at high cost).
@@
Whenever the price is “sufficiently attractive”, the pseudo-cost minimization finds the best decision that purchase just enough to maintain a certain competitive ratio. To define this trade-off, Lechowicz et al. (2024) introduce a dynamic threshold function $\phi(w):[0,1]\rightarrow[p_{\min},p_{\max}]$, which is a function that maps the current amount that has been purchased $w$ to a price value. It is defined as:
(8) $\phi(w)=p_{\max}-\gamma+\left(\frac{p_{\max}}{\alpha_{\texttt{RORO}}}-p_{\max}+2\gamma\right)\exp\left(\frac{w}{\alpha_{\texttt{RORO}}}\right),$
@@
where $\alpha_{\texttt{RORO}}$ is the optimal competitive ratio for OCS, defined as the unique positive solution to $\exp\left(\frac{1}{\alpha_{\texttt{RORO}}}\right)=\frac{p_{\max}-p_{\min}-2\gamma}{\nicefrac{{p_{\max}}}{{\alpha_{\texttt{RORO}}}}-p_{\max}-2\gamma}$ (Lechowicz et al., 2024):
@@
(9) $\alpha_{\texttt{RORO}}\coloneqq\left[W\left(\left(\frac{2\gamma+p_{\min}}{p_{\max}}-1\right)\exp\left(\frac{2\gamma}{p_{\max}}-1\right)\right)-\frac{2\gamma}{p_{\max}}+1\right]^{-1}.$
@@
In the above, $W(\cdot)$ is the Lambert $W$ function (Corless et al., 1996). Given this definition of $\alpha_{\texttt{RORO}}$, note that $\phi(\cdot)$ is monotonically decreasing on the interval $w\in[0,1]$ (i.e., non-increasing in the amount of demand that has already been purchased), and captures the marginal trade-off of making progress towards satisfying the single unit of demand. We summarize the RORO algorithm in Algorithm 3.
Algorithm 3 Online Ramp-On, Ramp-Off Algorithm (RORO(Lechowicz et al., 2024)